OpenStudy (anonymous):

Which of the following points does NOT lie on the graph of the following equation? 5x + 6y = -3 A.(-3, 2) B.(3, -3) D.(9, -8) C.(11, -9)

OpenStudy (anonymous):

C because 5(11) -54 = -1

OpenStudy (anonymous):

dude do it the same as the way I showed you lol 5(11)+9(-9)=10 10=/=-3 answer is c again.

OpenStudy (anonymous):

im just trying to get my summer school work done fast

OpenStudy (anonymous):

maybe you should just do your work bud..

OpenStudy (anonymous):

maybe im runing out of time

jhonyy9 (jhonyy9):

D. is right

jhonyy9 (jhonyy9):

45-48 = -3